# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v9

@@ 3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017). The CULane dataset (Pan et al., 2018) serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing. In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder (Harald, 1924) is represented as:
@@
To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details. The Depth-specific Attention (DSA) generates an attention map from the original image. Once processed through the DSA and STE Modules, as depicted in Fig. 5, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map. The Depth Prediction (DP) module is then responsible for estimating the depth.
@@
According to the transmittance formula Eq. (2), we use the dark channel prior(He et al., 2011) to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
@@ 4.1 Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods. As illustrated in the Fig. 8, the network comprises four main components. To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network. In the neck network, we utilize HRNetV2(Wang et al., 2021) to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
@@
For the detection head, we employ the concept of dynamic convolution, incorporating both lane instance detection and lane mask segmentation. During instance detection, the network identifies lane instances within the image and predicts a set of dynamic convolution kernel parameters for each lane. These parameters are subsequently applied to the lane mask segmentation head to filter the instance-specific segmentation map for each lane. This approach eliminates the need for complex post-processing steps, enhancing detection speed.
Moreover, to fully exploit the structural relationships between lane lines, we designed a kernel feature fusion module, which allows the network to automatically learn the correlations between predicted lane instances. This contributes to more accurate lane identification and localization. Additionally, to compensate for the loss of edge information in foggy images, we incorporated a low-level edge enhancement module, further improving the network’s ability to detect lane boundaries with precision.
@@ 4.2 Global Feature Fusion Module
In recent years, Transformer architectures have made considerable strides in the field of natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025; Chen et al., 2023). However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
@@
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers(Liu et al., 2021c, 2022; Dong et al., 2022). Among them, the Swin Transformer(Liu et al., 2021c) combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
@@ 4.3 Kernel Feature Fusion Module
Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution. CondInst(Tian et al., 2023) is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency. Similarly, CondLane(Liu et al., 2021a) applies dynamic convolution for lane instance segmentation, achieving strong performance.
Building on this, our lane detection head also uses dynamic convolution, consisting of lane instance detection and mask segmentation. Unlike traditional methods, we generate dynamic convolution kernels using the lane’s starting position features rather than center features, which are more suitable for the elongated nature of lanes. We design an instance detection head with two branches: one predicts lane starting points and generates a heatmap, and the other generates dynamic convolution kernels for lane segmentation.
However, existing methods don’t fully utilize the structural relationships between lane instances. To address this, we introduce a Kernel Feature Fusion Module (KFFM) that enables the network to learn and predict relationships between lane instances, enhancing detection performance. KFFM is shown in the bottom right of Fig. 8.
@@ 4.4 Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 10. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@
To address the challenges of lane detection in foggy conditions, we apply global feature fusion within both the backbone network and detection head to enhance high-dimensional global features. Additionally, we use multi-scale fusion to aggregate detail features from low-resolution feature maps. However, this approach alone does not fully compensate for the loss of edge detail inherent in foggy images. Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a 3 × 3 and a 1 × 1 convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 8. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
@@ 4.5.5 Edge prediction
For the task of edge enhancement, the method treats it as an edge segmentation problem, which is essentially a binary classification task. The labels at edge locations are assigned a value of 1, while non-edge locations are assigned 0. Accordingly, the cross-entropy loss function is employed for optimization. Let $e_{gt}^{i}$ denote the label of pixel $i$, and $p_{i}$ represent the predicted probability of the pixel being positive. The loss function is defined as:
@@ 5.1 Datasets
Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
[…]
  • 1.
@@ 5.2 Evaluation Metric
Lane detection is typically approached as a semantic segmentation problem, assessed using common metrics. However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics. Adhering to SCNN’s (Pan et al., 2018) framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions. Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct. From this, we calculate Precision, Recall, and F1-score using the following formulas:
@@
Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@ 5.3 Implementation Details
To benchmark against existing methods(Pan et al., 2018; Zheng et al., 2021; Tabelini et al., 2021a; Zheng et al., 2022; Ran et al., 2023; Liu et al., 2021b; Dai et al., 2024; Qin et al., 2020; Liu et al., 2021a; Wang et al., 2025), we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
@@ Acknowledgement
This project is jointly supported by National Natural Science Foundation of China (Nos. 52172350, W2421069 and 51775565), the Guangdong Basic and Applied Research Foundation (No. 2022B1515120072), the Guangzhou Science and Technology Plan Project (No. 2024B01W0079), the Nansha Key RD Program (No. 2022ZD014), the Science and Technology Planning Project of Guangdong Province (No. 2023B1212060029).We would like to release our source code and dataset.

### v10

@@ 3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane[8] and TuSimple[24]. The CULane dataset [8] serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing. In contrast, the TuSimple dataset [24] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model [18], and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder [44] is represented as:
@@
To calculate the corresponding $d(x)$, we use the S2R-DepthNet[45] that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details. The Depth-specific Attention (DSA) generates an attention map from the original image. Once processed through the DSA and STE Modules, as depicted in Fig. 5, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map. The Depth Prediction (DP) module is then responsible for estimating the depth.
@@
According to the transmittance formula Eq. (2), we use the dark channel prior[46] to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
@@ 4.1 Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network. The network works through several collaborative modules to tackle the challenges posed by foggy images, with the overall architecture shown in Fig. 8. The input image first passes through the backbone network, where feature extraction occurs, followed by the Global Feature Fusion Module (GFFM). This module is designed to mitigate the loss of global information in foggy images, allowing the network to retain essential global context. The output is a feature map enriched with both local and global information, crucial for accurate lane detection.
The multi-level feature map from the backbone is then passed into the Neck section of the network, which employs HRNetV2[47] for multi-scale feature fusion. HRNetV2 excels in integrating high-resolution and low-resolution features from different stages of the backbone. This process helps the network capture spatial information at multiple scales, producing a rich, multi-scale feature representation that forms the basis for lane instance detection.
Next, the Instance Detection Head processes the multi-scale feature map from the Neck network, predicting dynamic convolution kernel parameters for each lane instance. These dynamic kernels are applied to the lane mask segmentation head to isolate instance-specific features. The output includes a parameter map and a heat map, crucial for identifying lane instances. This approach eliminates the need for complex post-processing, significantly enhancing detection speed and accuracy.
The output parameter map is then passed to the Kernel Feature Fusion Module (KFFM), which learns the relationships between different lane instances. The KFFM fuses these instance-specific features, improving lane localization accuracy, especially for complex or overlapping lane lines. This module ensures that the network can better identify lane instances and refine the localization of each lane.
Finally, to compensate for the loss of edge information in foggy images, the network incorporates a Low-Level Edge Enhancement Module (LEEM). This module enhances edge features using convolution operations, recovering fine details at lane boundaries. The enhanced edge features improve the network’s ability to detect lane boundaries with high precision, even under challenging foggy conditions.
@@ 4.2 Global Feature Fusion Module
As shown in Fig. 9, to enhance global feature integration in the Swin Transformer backbone[48], we replace the final window-based attention mechanism with a standard multi-head attention mechanism (MHSA) from Vision Transformer (ViT)[49]. While ViT excels at capturing global dependencies, it suffers from high computational cost, especially at high resolutions, making it unsuitable for real-time tasks such as lane detection. The traditional ViT computes global self-attention with a complexity of $O(N^{2})$, where $N$ is the number of input tokens. This high computational cost becomes prohibitive at high resolution, slowing down inference and limiting real-time processing.
In our approach, we address this issue by introducing an efficient solution: global attention is applied only after reducing the image resolution. This ensures that the computational cost of global attention is manageable, allowing us to retain the global attention mechanism of ViT while maintaining fast inference speed, which is crucial for real-time lane detection in foggy conditions.
In our design, this MHSA layer replaces the original window-based attention mechanism used in the Swin Transformer. By applying global attention after reducing the image resolution, we strike a balance between maintaining global context awareness and achieving high computational efficiency, which is crucial for real-time lane detection tasks.
@@ 4.3 Kernel Feature Fusion Module
Existing lane instance segmentation methods[50, 36], typically use dynamic convolution to generate kernel parameters for each instance, which improves the network’s flexibility and adaptability. However, these methods do not fully leverage the structural relationships between lane instances, leading to limited segmentation and detection accuracy. Traditional dynamic convolution methods only consider individual lane instances and overlook the potential relationships between adjacent instances, which reduces their performance when handling complex lane scenes.
To address this limitation, we propose the Kernel Feature Fusion Module (KFFM), as shown in the bottom right of Fig. 8, which optimizes kernel generation by integrating information from adjacent lane instances. Specifically, KFFM introduces structured feature fusion into dynamic convolution, allowing each lane instance’s kernel to depend not only on the features of the current lane but also on the information from nearby lane instances. This approach enables the network to adaptively learn the relationships between lane instances, improving both detection accuracy and segmentation performance.
Let the input feature map from the lane instance detection branch be denoted as $\mathit{K}=\left[\mathit{kernel}_{0},\mathit{kernel}_{1},\dots,\mathit{kernel}_{N}\right]$, where each $\mathit{kernel}_{i}\in\mathbb{R}^{k\times k\times C}$, and define the weight matrix $\mathit{W}\in\mathbb{R}^{(N+1)\times(N+1)}$ to capture the contributions of each kernel to the final fused kernel.
The final fused kernel for lane $i$ is computed as the weighted sum of the kernels:
$\hat{\mathit{kernel}}_{i}=\sum_{j=0}^{N}W_{i,j}\cdot\mathit{kernel}_{j}$ (4)
where $W_{i,j}$ is the weight for the contribution of $\mathit{kernel}_{j}$ to the fused kernel $\hat{\mathit{kernel}}_{i}$; The weight matrix $\mathit{W}$ is learned during training.
The final fused kernel is obtained through matrix multiplication as follows:
$\hat{\mathit{K}}_{i}=\mathit{W}_{i}\cdot\mathit{K}$ (5)
where $\mathit{W}_{i}$ corresponds to the row of the weight matrix $\mathit{W}$ associated with kernel $i$.
The resulting fused kernel $\hat{\mathit{kernel}}_{i}$ is then used in the dynamic convolution operation for lane segmentation, leading to more accurate lane detection and segmentation.
@@ 4.4 Low-level Edge Enhanced Module
Images captured in foggy conditions differ from those captured in good lighting primarily due to blurriness and low contrast, which makes the edges of objects and scene details indistinct and difficult to detect. This degradation of edge information significantly impacts the algorithm’s ability to accurately understand and recognize the scene. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial relationships about objects, which is critical for tasks such as object detection, segmentation, and tracking. We used the Canny edge detection algorithm with identical threshold settings to process both foggy and normal images, as shown in Fig. 10. The results show that fog significantly degrades edge information, particularly lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—accurate edge detection is essential for reliable lane detection.
To address the challenges posed by foggy conditions, we propose the integration of global feature fusion within both the backbone network and detection head to enhance high-dimensional global features. Additionally, we incorporate multi-scale fusion to aggregate detailed features from low-resolution feature maps. However, this approach alone does not fully compensate for the loss of edge details in foggy images. To resolve this issue, we introduce the Low-level Edge Enhanced Module (LEEM) into the network. LEEM focuses on edge information, effectively recovering and refining the edges that are crucial for lane detection, thus improving both the accuracy and stability of lane detection in foggy environments.
@@
LEEM works by adding an auxiliary path to the network that specifically learns edge features of the input image. The process begins with the application of the Canny operator for edge detection, followed by the reconstruction of missing lane edges using lane line labels. This ground truth for the edge enhancement module is generated by combining detected edges with labeled edges. This approach helps the network better identify lane boundaries, which are often difficult to distinguish in foggy conditions.
In the second stage of the backbone network, we introduce a branch that processes the extracted edge features. This branch sequentially applies a $3\times 3$ convolution followed by a $1\times 1$ convolution to enhance edge-related features. The mathematical representation of this process is as follows:
$\hat{\mathit{E}}_{\mathit{predict}}=\mathit{Conv}_{2}\left(\mathit{ReLU}\left(\mathit{BN}(\mathit{Conv}_{1}(\mathit{X})))\right)\right)$ (6)
where $\mathit{Conv}_{1}$ is a $3\times 3$ convolution designed to capture spatial edge features, $\mathit{BN}$ is the batch normalization layer to stabilize the learning process by normalizing the output, $\mathit{ReLU}$ is the activation function that provides non-linearity, $\mathit{Conv}_{2}$ is a $1\times 1$ convolution used to reduce the dimensionality and refine edge features.
During training, the edge information extracted via the Canny operator from both the input image and ground truth (GT) lane lines is first downsampled to match the resolution of the network’s predicted edge information $\hat{\mathit{E}}_{\mathit{predict}}$. Then, these two edge information maps are merged. The final combined edge information is represented as $\hat{\mathit{E}}_{\mathit{GT}}$, which can be computed as:
$\hat{\mathit{E}}_{\mathit{GT}}=\mathit{Edge}_{I}+\mathit{Edge}_{\mathit{Lane}}$ (7)
where $\mathit{Edge}_{I}$ is the edge information obtained from the input image $\mathit{I}$ using the Canny operator, $\mathit{Edge}_{\mathit{Lane}}$ is the edge information derived from the lane ground truth (GT) using the Canny operator.
Next, we compute the cross-entropy loss between the predicted edge information $\hat{\mathit{E}}_{\mathit{predict}}$ and the final combined edge information $\hat{\mathit{E}}_{\mathit{GT}}$. This loss function is formulated as:
$L_{\mathit{edge}}=-\sum_{i}\left(e_{\mathit{gt}}^{i}\log(p_{i})+(1-e_{\mathit{gt}}^{i})\log(1-p_{i})\right)$ (8)
where $e_{\mathit{gt}}^{i}$ represents the ground truth label for pixel $i$ (1 if it’s part of the edge, 0 otherwise), $p_{i}$ is the predicted probability that pixel $i$ is part of the edge.
The LEEM module allows the network to effectively enhance critical edge details in lane detection, significantly improving both the accuracy and stability of lane detection in foggy conditions. This module not only recovers lost edge information in images but also integrates information from various sources, using the cross-entropy loss function to optimize the network. As a result, LEEM plays a key role in enhancing the robustness of lane detection systems in challenging environments.
@@
Figure 11: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
[…]
  • 1.
@@ 4.5.5 Edge prediction
For the task of edge enhancement, the method treats it as an edge segmentation problem, which is essentially a binary classification task. The labels at edge locations are assigned a value of 1, while non-edge locations are assigned 0. Accordingly, the cross-entropy loss function is employed for optimization.
As discussed in Section 4.4, the loss function for edge prediction is defined as:
@@
where $e_{gt}^{i}$ denotes the label of pixel $i$, and $p_{i}$ represents the predicted probability of the pixel being part of the edge.
@@ 5.2 Evaluation Metric
Lane detection is typically approached as a semantic segmentation problem, assessed using common metrics. However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics. Adhering to SCNN’s [8] framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions. Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct. From this, we calculate Precision, Recall, and F1-score using the following formulas:
@@
Figure 12: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@ 5.3 Implementation Details
To benchmark against existing methods[8, 35, 36, 25, 31, 28, 16, 32, 29, 51], we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
@@ Acknowledgement
This project is jointly supported by National Natural Science Foundation of China (Nos. 52172350, W2421069 and 51775565), the Guangdong Basic and Applied Research Foundation (No. 2022B1515120072), the Guangzhou Science and Technology Plan Project (No. 2024B01W0079), the Science and Technology Planning Project of Guangdong Province (No. 2023B1212060029).We would like to release our source code and dataset.